The susceptibility of annealed iron at saturation is 5500. Find the permeability of annealed iron at saturation.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Permiability (μ) = μ0(1 + x) Given susceptibility = 5500 μ = 4 × 10^–7 (1 + 5500) = 4 × 3.14 × 10^–7 × 5501 6909.56 × 10^–7 ≈ 6.9 × 10^–3